# Enhancing the De-identification of Personally Identifiable Information in Educational Data

### v1

@@ III-B Models for PII Detection
User Determine if {entity} is a privately identifiable information in its context: {context}, think carefully before saying no to protect against PII leakage, only output T or F. Determine if {entity} is a privately identifiable information in its context: {context}. Think step-by-step before outputting T or F.
Assistant “T” or “F” {CoT Reasoning} + T” or “{CoT Reasoning} + F
Verifier Models. To improve the precision (Equation 1) of PII detection while maintaining recall (Equation 2), we propose the addition of a verifier model as a second step to verify whether predictions are accurately identified as PII. This approach is inspired by recent studies [34, 35] that integrate verifiers into multi-step reasoning tasks in LLMs. For the implementation of verifier models, we propose two variants: Verifier Model I (Without CoT) and Verifier Model II (With CoT). These verifier models assess detected entities within their surrounding context to eliminate false positives while retaining true PII entities, thereby enhancing the precision of PII detection systems. Both verifiers are fine-tuned versions of the GPT-4o-mini model. The process of the verifier model approach is illustrated in Fig. 1 (⑤).
The Verifier Model II (With CoT) incorporates Chain-of-Thought (CoT) reasoning, a method shown to enhance decision-making by breaking down complex problems into intermediate steps [36]. The Verifier Model II (With CoT) aims to improve interpretability and robustness by generating reasoning before the final classification. However, it generates more output tokens, leading to higher computational costs. The Verifier Model I (Without CoT), which directly classifies entities without reasoning, is retained for scenarios where computational resources are limited or speed is a higher priority.
A subset of the CRAPII dataset is used to construct the training data for the verifier models. Detected entities are obtained from a chosen base model, and their context is extracted. Entities are labeled as T (True PII) or F (False PII) based on ground-truth labels in the CRAPII dataset. For the Verifier Model II (With CoT), we use GPT-4o-mini to generate reasoning to support the classification. If the reasoning does not align with the ground truth after six attempts, the label defaults to T to avoid mistakenly removing true PII entities and prioritize privacy preservation. The prompt structures for both verifier models are shown in Table V.
@@
III-C Splitting Data for Training and Testing
Our study focuses on four specific categories from the CRAPII dataset for PII detection: NAME, URL_PERSONAL, EMAIL, and PHONE_NUM. To support the training and evaluation requirements of our experiments, the 22,688 files in the CRAPII dataset are split into three distinct sets: Base Train Set (25%, 5,672 files), Verifier Train Set (15%, 3,403 files), and Test Set (60%, 13,613 files). This split ensures that all sets contain a sufficient number of entities from each category, including rare categories such as PHONE_NUM, which only has 15 entities in total. Table VI presents the distribution of true entity counts across the three sets.
TABLE VI: True Entity Counts Across Data Splits
@@ III-C Splitting Data for Training and Testing
III-D Evaluation Metrics
When evaluating the performance of a model designed to detect PII, it is essential to assess the model’s ability to correctly identify true PII entities while minimizing incorrect classifications. The True Positives (TP) indicate the number of correctly identified PII entities. The False Positives (FP) represent the number of non-PII entities incorrectly classified as PII, which can reduce the utility of the dataset by unnecessarily removing valuable information. The False Negatives (FN) denote the number of missed detections of actual PII entities, which could result in privacy breaches and violations of legal regulations. To evaluate the model’s capability in accurately detecting PII, we adopt the evaluation method suggested in [15]. This method considers a PII entity to be correctly identified only if it is an exact match with the corresponding entity in the text data. To provide a comprehensive assessment of the model’s performance, we consider multiple metrics that capture different aspects of the model’s effectiveness.
@@ III-D Evaluation Metrics
III-E Analysis of Cultural and Gender Bias in Name Detection by Models
@@
Evaluating the model’s performance in detecting <NAME_STUDENT> entities across different cultural and gender distributions is quite important. Models trained on imbalanced datasets may underperform in identifying names of groups that are underrepresented in training data. Specifically, if the training data lacks sufficient representation of names from certain cultural backgrounds, the model may exhibit a lower recall or precision for those names. This dimension of model evaluation is necessary to ensure fairness and inclusivity in PII detection systems, especially as names serve as direct identifiers with critical privacy implications.
@@
We analyzed the cultural and gender distributions of the names in the <NAME_STUDENT> entities in detail by using a two-step approach. First, we used a rule-based name parser [^10] https://github.com/derek73/python-nameparser to split each name into components, typically a first name and a last name. We then determined the gender of each name based on the first name and matched the nationality using the last name. This method aligns with the approach described in the CRAPII paper [14]. In the second step, we mapped the identified countries to their respective regional cultures using the ISO-3166 dataset with the UN regional codes [^11] https://github.com/lukes/ISO-3166-Countries-with-Regional-Codes. Specifically, we relied on the all.format file, which includes detailed regional and sub-regional classifications for each country. For example, “Nigeria” maps to the Africa region, while “United States of America” maps to the Americas. The ISO-3166 dataset provides five cultural regions: Asia, Americas, Europe, Africa, and Oceania. Notably, no names in the CRAPII dataset belonged to Oceania, so we focused on the remaining four cultural groups.
III-F Analysis of Models’ Generalizability
@@
To analyze the generalizability of our investigated models, we employed the TSCC dataset, as introduced in Section III-A. Notably, the TSCC dataset has been pre-redacted, with most redacted words replaced by placeholders indicating their name entity types, such as <STUDENT> and <TEACHER>. To evaluate the model’s ability to generalize across diverse contexts, it was necessary to replace these placeholders with synthetic entities that reflect diversity in gender and cultural backgrounds. This replacement ensures a realistic evaluation of the models’ performance when encountering unseen data with varying demographic characteristics. To generate a diverse and representative set of synthetic names, we systematically created a mapping of first and last names categorized by gender and cultural groups. This process involved the following steps:
@@
    For each culture, we identified its corresponding countries based on the all.format file in the United Nations dataset (all.csv [^12] https://github.com/lukes/ISO-3166-Countries-with-Regional-Codes/blob/master/all/all.csv).
@@
    For each country, we used the get_top_names function from the names-dataset Python package [^13] https://github.com/philipperemy/name-dataset to retrieve first names based on gender and last names based on the country, as outlined in the CRAPII paper [14].
@@
    We combined the names of all the countries that belong to that culture into one list, filtering out names containing non-English characters to maintain consistency with the TSCC dataset [10].
@@
This process resulted in a realistic and culturally diverse dataset that retains the original conversational structure while introducing diversity in entity representation. For instance, the processed version of the transcript shown in Section III-A appears as follows:
@@
IV Results and Discussion
TABLE VII: Performance Metrics for Different PII Detection Models
@@ 5. Fine-tuned
The performance metrics of all proposed PII detection models are summarized in Table VII.
IV-A Overall-Level Analysis of Model Performance
IV-A1 Presidio Models
For the two Presidio models utilizing different SpaCy configurations, en_core_web_trf consistently outperforms en_core_web_lg across all metrics, as shown in Table VII. The transformer-based en_core_web_trf achieves higher overall precision (0.2092 vs. 0.1505) and recall (0.8368 vs. 0.7103), likely due to its enhanced ability to capture long-range dependencies in text. However, both configurations exhibit low precision, which is likely due to the inclusive detection approach of the models. Although this approach enables the identification of a wide range of entities, including less common ones, it also leads to a significant number of false positives, thereby reducing overall precision.
IV-A2 Azure AI Language
@@ IV-A3 Prompting GPT-4o-mini
IV-A4 Fine-tuned GPT-4o-mini
@@
IV-A5 Verifier Models
The Verifier Model I (Without CoT) achieves the highest precision (0.8893) among all models by not defaulting to retaining entities when uncertain. However, this less conservative approach leads to the wrong removal of some true positives during verification, resulting in a lower recall of 0.8023. The Verifier Model II (With CoT) demonstrates a more balanced trade-off with a recall of 0.8648, as its conservative behavior defaults to retaining entities (T) when uncertainty arises. Notably, the precision scores for both verifier models surpass that of all other five methods, aligning with our effort to improve precision. However, their recall is lower than that of the Fine-tuned GPT-4o-mini model, suggesting that these verifier models may be better suited for tasks that prioritize precision over exhaustive detection.
IV-B PII Category-Level Analysis of Model Performance
IV-B1 Name Detection (NAME_STUDENT)
@@
IV-B2 URL Detection (URL_PERSONAL)
The Fine-tuned GPT-4o-mini model achieves the highest recall (0.9387) and $F_{5}$ score (0.9069), indicating a strong performance in capturing true positives. However, its precision (0.4914) remains moderate, suggesting room for improvement in reducing false positives. The Verifier Model I (Without CoT) significantly improves precision, achieving the highest value (0.9877) and $F_{1}$ score (0.8587), but at the cost of reduced recall (0.7594). This demonstrates the verifier model’s effectiveness in filtering false positives while highlighting the inherent trade-off between precision and recall, as no model achieves high performance in both metrics simultaneously.
IV-B3 Email Detection (EMAIL)
@@
IV-B4 Phone Number Detection (PHONE_NUM)
@@
IV-C Impact of Low-Precision PII Detection: Examples of Semantic Disruption
@@
The incorrect anonymization of names changes the intended meaning of the text, as the connection between “Nazareth” and its historical and religious significance is lost. This disruption affects students’ understanding and the utility of the data in educational contexts. Moreover, anonymization negatively impacts the utility of the text for machine learning applications. If the anonymized text is used to train or fine-tune a language model or included in a knowledge base for a Retrieval-Augmented Generation (RAG) pipeline, it may result in incorrect or misleading responses to queries about “Nazareth” [37]. Repeated inclusion of anonymized instances, such as associating “Nazareth” with unrelated names like “Elias Carson,” may erode the model’s understanding of its cultural significance, leading to inaccuracies in subsequent applications [38, 39].
@@
Replaced: Consider hiring a copywriter to craft a compelling https://techwaveinsight.ioeep menus clean – no grease and no food or water stains. Get rid of worn or torn menus.Update menu and prices at least once a http://elitecodingacademy.org menu around popular items.
@@
IV-D Cost Analysis
TABLE VIII: Cost Breakdown for PII Detection Models (in USD)
@@ IV-D Cost Analysis
Table VIII presents the cost associated with each PII detection model, complementing the performance metrics in Table VII. The results emphasize the potential of GPT-based approaches, particularly the Fine-tuned GPT-4o-mini model, for high-quality PII detection at significantly lower costs.
@@
IV-E Name Culture and Gender Bias Analysis
Based on the detection performance in Table VII, we selected three models—Presidio with en_core_web_trf, Azure AI Language, and Fine-tuned GPT-4o-mini—for the analysis of cultural and gender bias in name detection. The total number of entities for each gender and culture group and their recall are shown in Table IX.
TABLE IX: Recall comparison across gender and culture groups for selected models.
@@ IV-E Name Culture and Gender Bias Analysis
Culture Analysis. The cultural analysis reveals more significant differences in model performance. Both Microsoft models show lower recall for African and Asian names. In particular, Presidio exhibits a notable performance gap, with recall rates of 0.7647 for African names and 0.8640 for Asian names, compared to 0.9024 for European and 0.9091 for American names. Azure AI Language also shows a lower recall for African names (0.9244), although the gap is less pronounced than in Presidio. These results suggest inherent biases in the Microsoft models, possibly stemming from imbalances in the training data or gaps in cultural representation. In contrast, the fine-tuned GPT-4o-mini model achieves consistent and high recall across all cultural groups. It performs with recall scores of 0.9756 for European names, 0.9790 for American names, 0.9840 for Asian names, and 0.9748 for African names. This minimal variation in recall across cultural groups demonstrates the model’s ability to mitigate cultural bias and generalize effectively across diverse name distributions.
@@
IV-F Generalizability Analysis
TABLE X: Performance Metrics for Different PII Detection Approaches on the TSCC Dataset.

### v2

@@ 3.2 Models for PII Detection
User Determine if {entity} is a privately identifiable information in its context: {context}, think carefully before saying no to protect against PII leakage, only output T or F. Determine if {entity} is a privately identifiable information in its context: {context}. Think step-by-step before outputting T or F, format your response as (your reasoning) + [Response:] T or F
Assistant ‘‘T’’ or ‘‘F’’ ‘‘{CoT Reasoning} + T’’ or ‘‘{CoT Reasoning} + F’’
@@
GPT-4o-mini Training and Inference. We set temperature = 0 and top_p = 1 for both our GPT-4o-mini-based PII detector models (prompted and fine-tuned) and all verifier models to ensure strict adherence to the required output formats as suggested in the prior research [Wang et al. (2025)]. For the fine-tuning process, we adopted the hyperparameters recommended by the OpenAI platform: epochs = 2, batch_size = 1, and a learning_rate_multiplier = 1.8. No further hyperparameter tuning was performed due to these platform constraints.
3.3 Splitting Data for Training and Testing
Our study focuses on four specific categories from the CRAPII dataset for PII detection: NAME, URL_PERSONAL, EMAIL, and PHONE_NUM. To support our multi-stage experimental design, which involves training base models and then separate verifier models, the 22,688 files in the CRAPII dataset are split into three distinct sets: Base Train Set (25%, 5,672 files), Verifier Train Set (15%, 3,403 files), and Test Set (60%, 13,613 files). This split ensures that all sets contain a sufficient number of entities from each category, including rare categories such as PHONE_NUM, which only has 15 entities in total. Table 6 presents the distribution of true entity counts across the three sets.
Table 6: True Entity Counts Across Data Splits
@@ 3.3 Splitting Data for Training and Testing
3.4 Evaluation Metrics
[…]
@@ 3.4 Evaluation Metrics
3.5 Analysis of Cultural and Gender Bias in Name Detection by Models
@@
Evaluating the model’s performance in detecting <NAME_STUDENT> entities across different cultural and gender distributions is quite important. We adopt Hofstede’s definition of culture as “the collective programming of the mind that distinguishes the members of one group or category of people from another[Hofstede (2001)]. Because this collective programming shapes online expression and data creation, cultural groups are not represented equally in large-scale datasets. Training data, often sourced from the internet, tends to reflect the disproportionate influence of dominant cultures, leading to their overrepresentation. A similar imbalance exists for gender, where available data often feature a skewed distribution of male and female names. Models trained on imbalanced datasets may underperform in identifying names of groups that are underrepresented in training data. Specifically, if the training data lacks sufficient representation of names from certain cultural backgrounds, the model may exhibit a lower recall or precision for those names. This dimension of model evaluation is necessary to ensure fairness and inclusivity in PII detection systems, especially as names serve as direct identifiers with critical privacy implications.
@@
We analyzed the cultural and gender distributions of the names in the <NAME_STUDENT> entities in detail by using a two-step approach. First, we used a rule-based name parser [^10] https://github.com/derek73/python-nameparser to split each name into components, typically a first name and a last name. We then determined the gender of each name based on the first name and matched the nationality using the last name. This method aligns with the approach described in the CRAPII paper [Holmes et al. (2024)]. In the second step, we mapped the identified countries to their respective regional cultures using the ISO-3166 dataset with the UN regional codes. [^11] https://github.com/lukes/ISO-3166-Countries-with-Regional-Codes Specifically, we relied on the all.format file, which includes detailed regional and sub-regional classifications for each country. For example, “Nigeria” maps to the Africa region, while “United States of America” maps to the Americas. The ISO-3166 dataset provides five cultural regions: Asia, Americas, Europe, Africa, and Oceania. Notably, no names in the CRAPII dataset belonged to Oceania, so we focused on the remaining four cultural groups.
To quantify potential biases, we evaluate the model for fairness using the principle of Equality of Opportunity [Hardt et al. (2016)]. This metric assesses whether the model correctly identifies student names at equal rates across different protected groups. We measure this by computing the true positive rate (TPR) for each cultural and gender subgroup, where fairness is achieved if the following condition holds:
$\forall{A},P[\hat{Y}=1|A=1,Y=1]=P[\hat{Y}=1|A=0,Y=1]$
Here, $Y=1$ represents a true <NAME_STUDENT> entity, and $A$ is an indicator variable for membership in a protected group (e.g., $A=1$ for ‘Female,’ $A=0$ for ‘Male’). This metric evaluates if the probability of a correct identification ($\hat{Y}=1$) is independent of group membership ($A$), given that a name is actually present ($Y=1$).
Our decision to aggregate countries into four continental cultural groups was driven by data sparsity. Although a country is a more granular cultural indicator, our data contains names from many countries, with most being represented by only a few samples. Therefore, we opted for this coarser grouping. This methodological choice is supported by recent research demonstrating a convergence of cultural values at the continental level [Jackson and Medvedev (2024)]. Similarly, our gender analysis is limited to a binary classification because, to our knowledge, there is no reliable computational method for mapping personal names to non-binary genders.
3.6 Analysis of Models’ Generalizability
@@
To analyze the generalizability of our investigated models, we employed the TSCC dataset, as introduced in Section 3.1. Notably, the TSCC dataset has been pre-redacted, with most redacted words replaced by placeholders indicating their name entity types, such as <STUDENT> and <TEACHER>. To evaluate the model’s ability to generalize across diverse contexts, it was necessary to replace these placeholders with synthetic entities that reflect diversity in gender and cultural backgrounds. This replacement ensures a realistic evaluation of the models’ performance when encountering unseen data with varying demographic characteristics. To generate a diverse and representative set of synthetic names, we systematically created a mapping of first and last names categorized by gender and cultural groups. This process involved the following steps:
@@
    For each culture, we identified its corresponding countries based on the all.format file in the United Nations dataset.
@@
    For each country, we used the get_top_names function from the names-
    dataset
    Python package [^12] https://github.com/philipperemy/name-dataset to retrieve first names based on gender and last names based on the country, as outlined in the CRAPII paper [Holmes et al. (2024)].
@@
    We combined the names of all the countries that belong to that culture into one list, filtering out names containing non-English characters to maintain consistency with the TSCC dataset [Caines et al. (2022)].
@@
This process resulted in a realistic and culturally diverse dataset that retains the original conversational structure while introducing diversity in entity representation. For instance, the processed version of the transcript shown in Section 3.1 appears as follows:
@@
4 Results
Table 7: Performance Metrics for Different PII Detection Models (The highest values across all models are shown in bold to indicate the best performance for each metric.)
@@ 5. Fine-tuned
The performance metrics of all proposed PII detection models are summarized in Table 7.
4.1 Overall-Level Analysis of Model Performance
4.1.1 Presidio Models
For the two Presidio models utilizing different spaCy configurations, en_core_web_trf consistently outperforms en_core_web_lg across all metrics, as shown in Table 7. The transformer-based en_core_web_trf achieves higher overall precision (0.2092 vs. 0.1505) and recall (0.8368 vs. 0.7103), likely due to its enhanced ability to capture long-range dependencies in text. However, both configurations exhibit low precision, which is likely due to the inclusive detection approach of the models. Although this approach enables the identification of a wide range of entities, including less common ones, it also leads to a significant number of false positives, thereby reducing overall precision.
4.1.2 Azure AI Language
@@ 4.1.3 Prompting GPT-4o-mini
4.1.4 Fine-tuned GPT-4o-mini
@@
4.1.5 Verifier Models
The Verifier Model I (Without CoT) achieves the highest precision (0.8893) among all models by not defaulting to retaining entities when uncertain. However, this less conservative approach leads to the wrong removal of some true positives during verification, resulting in a lower recall of 0.8023. The Verifier Model II (With CoT) demonstrates a more balanced trade-off with a recall of 0.8648, as its conservative behavior defaults to retaining entities when uncertainty arises. Notably, the precision scores for both verifier models surpass that of all other five methods, aligning with our effort to improve precision. However, their recall is lower than that of the Fine-tuned GPT-4o-mini model, suggesting that these verifier models may be better suited for tasks that prioritize precision over exhaustive detection.
4.2 PII Category-Level Analysis of Model Performance
4.2.1 Name Detection (NAME_STUDENT)
@@
4.2.2 URL Detection (URL_PERSONAL)
For the task of detecting personal URLs (links that can reveal personal information), the Fine-tuned GPT-4o-mini model achieves the highest recall (0.9387) and $F_{5}$ score (0.9069), indicating a strong performance in capturing true positives. However, its precision (0.4914) remains moderate, suggesting room for improvement in reducing false positives. The Verifier Model I (Without CoT) significantly improves precision, achieving the highest value (0.9877) and $F_{1}$ score (0.8587), but at the cost of reduced recall (0.7594). This demonstrates the verifier model’s effectiveness in filtering false positives while highlighting the inherent trade-off between precision and recall, as no model achieves high performance in both metrics simultaneously.
4.2.3 Email Detection (EMAIL)
@@
4.2.4 Phone Number Detection (PHONE_NUM)
@@
4.3 Impact of Low-Precision PII Detection: Examples of Semantic Disruption
@@
The incorrect anonymization of names changes the intended meaning of the text, as the connection between “Nazareth” and its historical and religious significance is lost. This disruption affects students’ understanding and the utility of the data in educational contexts. Moreover, anonymization negatively impacts the utility of the text for machine learning applications. If the anonymized text is used to train or fine-tune a language model or included in a knowledge base for a Retrieval-Augmented Generation (RAG) pipeline, it may result in incorrect or misleading responses to queries about “Nazareth” [Wang et al. (2025)]. Repeated inclusion of anonymized instances, such as associating “Nazareth” with unrelated names like “Elias Carson,” may erode the model’s understanding of its cultural significance, leading to inaccuracies in subsequent applications [Bender et al. (2021), Sun et al. (2020)].
@@
Replaced: Consider hiring a copywriter to craft a compelling
https://techwaveinsight.ioeep menus clean – no grease and no food or water stains. Get rid of worn or torn menus.Update menu and prices at least once a http://elitecodingacademy.org menu around popular items.
@@
4.4 Cost Analysis
Table 8: Cost Breakdown for PII Detection Models (in USD)
@@ 4.4 Cost Analysis
Table 8 presents the cost associated with each PII detection model, complementing the performance metrics in Table 7. The results emphasize the potential of GPT-based approaches, particularly the Fine-tuned GPT-4o-mini model, for high-quality PII detection at significantly lower costs.
@@
4.5 Name Culture and Gender Bias Analysis
Based on the results in Table 7, we selected three models—Presidio with en_core_web_trf, Azure AI Language, and Fine-tuned GPT-4o-mini—for the analysis of cultural and gender bias in name detection. To statistically validate these observations, we also perform two-tailed Mann-Whitney U tests. Each test evaluates whether a model’s performance on a specific protected group is significantly different from its performance on the aggregate of the other groups within the same category (e.g., recall on African names vs. all non-African names). The total number of entities for each gender and culture group and their recall are shown in Table 9. The results of these tests are shown in Table 10.
Table 9: Recall comparison across gender and culture groups for selected models
@@ 4.5 Name Culture and Gender Bias Analysis
Table 10: Statistics on TPR differences across gender and cultural groups for selected models, as measured by the Mann-Whitney U test
Category Group Model Group TPR Rest TPR U-statistic p-value
Gender Male Presidio 0.8786 0.8832 788937 n.s.
Azure 0.9494 0.9541 788888 n.s.
GPT-4o-mini 0.9646 0.9591 796957 n.s.
Female Presidio 0.8832 0.8786 796431 n.s.
Azure 0.9541 0.9494 796480 n.s.
GPT-4o-mini 0.9591 0.9646 788411 n.s.
Culture Europe Presidio 0.9024 0.8734 336670 n.s.
Azure 0.9585 0.9543 328579 n.s.
GPT-4o-mini 0.9756 0.9799 325760 n.s.
Americas Presidio 0.9091 0.8571 518076 ***
Azure 0.9755 0.9399 510039 ***
GPT-4o-mini 0.9790 0.9791 492456 n.s.
Asia Presidio 0.8640 0.8845 368796 n.s.
Azure 0.9320 0.9628 364898 ***
GPT-4o-mini 0.9840 0.9774 378976 n.s.
Africa Presidio 0.7647 0.8948 183022 ***
Azure 0.9244 0.9593 203048 **
GPT-4o-mini 0.9748 0.9796 209372 n.s.
Significance levels: n.s. = not significant, * $p<0.1$, ** $p<0.05$, *** $p<0.01$
@@
Culture Analysis. The cultural analysis reveals more significant differences in model performance. Both Microsoft models show significantly lower recall for African and Asian names. In particular, Presidio exhibits a notable performance gap, with recall rates of 0.7647 for African names and 0.8640 for Asian names, compared to 0.9024 for European and 0.9091 for American names. Azure AI Language also shows a lower recall for African names (0.9244), although the gap is less pronounced than in Presidio. These results suggest inherent biases in the Microsoft models, possibly stemming from imbalances in the training data or gaps in cultural representation. In contrast, the Fine-tuned GPT-4o-mini model achieves consistent and high recall across all cultural groups, demonstrating no statistically significant performance differences among the cultural groups. It performs with recall scores of 0.9756 for European names, 0.9790 for American names, 0.9840 for Asian names, and 0.9748 for African names. This minimal variation in recall across cultural groups demonstrates the model’s ability to mitigate cultural bias and generalize effectively across diverse name distributions.
@@
4.6 Generalizability Analysis
Table 11: Performance Metrics for Different PII Detection Approaches on the TSCC Dataset (The highest values across all models are shown in bold to indicate the best performance for each metric.)